# High resolution quantum enhanced phase imaging of cells

### v1

@@ III.2 Characterization of Twin-Beams by NRF
[…]
@@ III.3 TIE with phase and amplitude
Figure 6: Photon counts with the object placed at three distances. (Left panel) Photon counts recorded with the object positioned at the focal plane of the imaging system. In this case, only the transmittance mask ($\varnothing$) is visible. When the object is moved to defocused planes at distances ±$\pm dz$ (middle and right panel), phase variations induce additional intensity modulations, resulting in an overlap of phase and amplitude effects. Nevertheless, the TIE allows decoupling of these contributions, enabling the quantitative and unique retrieval of the sample’s phase profile.
In Fig. 6, the single-frame photon count images $N^{\prime}_{s}(\mathbf{x},z)$ are shown for three different axial positions of the object relative to the CFF plane: $z=0$ (left), $z=+dz$ (center), and $z=-dz$ (right) (with $dz=0.05$ mm). When the object is located at the nominal focus ($z=0$), only the amplitude (non-uniform transmittance) features are visible in the intensity distribution. In contrast, at defocused positions (±$\pm dz$), phase-induced propagation effects become evident, revealing the $\pi$-shaped phase structure superimposed on the amplitude modulation of the $\varnothing$-sign and its phase. Notably, the out-of-focus images exhibit overlapping features from both the amplitude and phase masks, underscoring the need for proper decoupling via the TIE to retrieve the pure phase profile.
A way to solve the TIE in Eq. (2) with the measured photon counts is to adopt the so-called Teague’s assumption Teague (1983), introducing an auxiliary function $\psi(\mathbf{x})$ so that
@@
According to Helmholtz’s theorem, the vector field $I(\mathbf{x},0)\nabla\phi(\mathbf{x},0)$ can be decomposed into an irrotational component and a solenoidal component: $I(\mathbf{x},0)\nabla\phi(\mathbf{x},0)=\nabla\psi+\nabla\times\mathbf{A}$, where $\psi$ is a scalar potential and $\mathbf{A}$ is a vector potential. It follows that Teague’s assumption holds if the transverse flux $I(\mathbf{x},0)\nabla\phi(\mathbf{x},0)$ is irrotational. It has been shown that the irrotational approximation is valid if the in-focus intensity distribution is nearly uniform Zuo et al. (2020). In this case, the phase discrepancy resulting from Teague’s auxiliary function is essentially negligible. However, when the measured sample exhibits low transmittance, the phase discrepancy may be relatively large and cannot be neglected. In our case, we consider a relatively high transmittance profile (see Fig. 1b), so we can safely use Teague’s assumption. The validity of the approximation is confirmed a posteriori by the correct recovery of the phase values, as presented in Fig. 2.
@@
$-k\frac{\partial}{\partial z}I(\mathbf{x},z)=\nabla^{2}_{\mathbf{x}}\psi(\mathbf{x}).$ (11)
@@ III.4.1 Phase resolution enhancement
A core aspect of this work is the demonstrated resolution enhancement, as shown in Fig. 3. This improvement stems from the suppression of high-frequency noise via the solution of the TIE, as described in Eq. (6). A visual explanation of this mechanism is provided in the inset images within the plot shown in Fig. 7.
Figure 7: Visualization of noise suppression in phase retrieval with $k_{\text{TIE}}$ estimator. The plot shows the percentage of shot-noise removed as a function of correlation length $l_{\text{CFF}}$. Insets show: (top) experimental Poissonian noise pattern $\sigma(\mathbf{x})$, (bottom left) its low-frequency component $\sigma_{\text{low}}(\mathbf{x})$, and (bottom right) the resulting phase noise $\phi_{\text{noise}}(\mathbf{x})$ after TIE processing.
@@ III.4.2 Phase resolution evaluation
To evaluate the spatial resolution of our imaging system, we analyze intensity profiles extracted along a line perpendicular to the edge of the phase samples averaged over $10^{3}$ frames, in order to eliminate any shot-noise. These intensity profiles are fitted using an error function (Edge Spread Function, ESF) centered at position $x_{0}$:
@@
To quantify the uncertainty, we first determine the 95% confidence interval for the fitted parameter $w$, indicated as $[w_{\text{sub}},w_{\text{sup}}]$. The standard error for $w$ is then approximated as $se_{w}=(w_{\text{sup}}-w_{\text{sub}})/3.92$ according to the $z$-test. Consequently, the standard error of the resolution $R$ is calculated as:
@@ V Author contributions
GO, IRB and AP devised the scheme of high resolution NIQPI. AP, GO and IRB developed the software for simulations and data analysis. The experiment and the data analysis have been carried out by AP with the help of IRB. IRB and AP wrote the paper with the contribution of all the Authors. The non-biological sample has been designed, fabricated and characterized by SS. IRB and MG are equally last co-Authors. IRB supervised the project together with MG who is the director of the Research Sector in which the work has been developed.

### v2

@@ III.2 Characterization of Twin-Beams by NRF
$\eta_{c}=\left(\frac{L_{\text{det}}}{M}\right)^{-2}\int_{\mathcal{A}}d\mathbf{x}_{s}\int_{\mathcal{A}}d\mathbf{x}_{i}\,\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(\mathbf{x}_{i}+\mathbf{x}_{s}+\Delta)^{2}}{2\sigma^{2}}}$ (8)
@@ III.3 TIE with phase and amplitude
Figure 5: Photon counts with the object placed at three distances. (Left panel) Photon counts recorded with the object positioned at the focal plane of the imaging system. In this case, only the transmittance mask ($\varnothing$) is visible. When the object is moved to defocused planes at distances $\pm dz$ (middle and right panel), phase variations induce additional intensity modulations, resulting in an overlap of phase and amplitude effects. Nevertheless, the TIE allows decoupling of these contributions, enabling the quantitative and unique retrieval of the sample’s phase profile.
In Fig. 5, the single-frame photon count images $N^{\prime}_{s}(\mathbf{x},z)$ are shown for three different axial positions of the object relative to the CFF plane: $z=0$ (left), $z=+dz$ (center), and z=dzz=-dz (right), with $dz=0.05$ mm. When the object is located at the nominal focus ($z=0$), only the amplitude (non-uniform transmittance) features are visible in the intensity distribution. In contrast, at defocused positions ($\pm dz$), phase-induced propagation effects become evident, revealing the $\pi$-shaped phase structure superimposed on the amplitude modulation of the $\varnothing$-sign and its phase. Notably, the out-of-focus images exhibit overlapping features from both the amplitude and phase masks, underscoring the need for proper decoupling via the TIE to retrieve the pure phase profile.
A way to solve the TIE in Eq. (2) with the measured photon counts is to adopt the so-called Teague’s assumption [61], introducing an auxiliary function $\psi(\mathbf{x})$ so that
@@
According to Helmholtz’s theorem, the vector field $I(\mathbf{x},0)\nabla\phi(\mathbf{x},0)$ can be decomposed into an irrotational component and a solenoidal component: $I(\mathbf{x},0)\nabla\phi(\mathbf{x},0)=\nabla\psi+\nabla\times\mathbf{A}$, where $\psi$ is a scalar potential and $\mathbf{A}$ is a vector potential. It follows that Teague’s assumption holds if the transverse flux $I(\mathbf{x},0)\nabla\phi(\mathbf{x},0)$ is irrotational. It has been shown that the irrotational approximation is valid if the in-focus intensity distribution is nearly uniform [72]. In this case, the phase discrepancy resulting from Teague’s auxiliary function is essentially negligible. However, when the measured sample exhibits low transmittance, the phase discrepancy may be relatively large and cannot be neglected. In our case, we consider a relatively high transmittance profile (see Fig. 1b), so we can safely use Teague’s assumption. The validity of the approximation is confirmed a posteriori by the correct recovery of the phase values, as reported in the caption of Fig. 2.
@@
$k\frac{\partial}{\partial z}I(\mathbf{x},z)=\nabla^{2}_{\mathbf{x}}\psi(\mathbf{x}).$ (11)
@@ III.4.1 Phase resolution enhancement
A core aspect of this work is the demonstrated resolution enhancement. This improvement stems from the suppression of high-frequency noise via the solution of the TIE, as described in Eq. (6). A visual explanation of this mechanism is provided in the inset images within the plot shown in Fig. 6.
Figure 6: Visualization of noise suppression in phase retrieval with $k^{(\text{TIE})}_{\text{opt}}$ estimator. The plot shows the percentage of shot-noise removed as a function of correlation length $l_{\text{CFF}}$. Insets show: (top) experimental Poissonian noise pattern $\sigma(\mathbf{x})$, (bottom left) its low-frequency component $\sigma_{\text{low}}(\mathbf{x})$, and (bottom right) the resulting phase noise $\phi_{\text{noise}}(\mathbf{x})$ after TIE processing.
@@ III.4.2 Phase resolution evaluation
Figure 7: Phase resolution evaluation with the optimized correction factor $k^{(\text{TIE})}_{\text{opt}}$ (a) Experimental and simulated phase resolution as a function of the resolution factor $\mathcal{D}$ for different defocus distances $dz$. The points correspond to experimental data for various dzdz values, while shaded areas represent theoretical predictions. As $\mathcal{D}$ increases, the resolution worsens due to the larger effective pixel size, while smaller $dz$ values allow better resolution. (b) Representative phase reconstructions for two different defocus distances ($dz=0.025$ mm and $dz=0.1$ mm) and three different $\mathcal{D}$ values. Insets show magnified regions of the sample for a better qualitative evaluation of the phase resolution changes.
Fig. 7 presents a quantitative evaluation of the spatial phase resolution achieved with the new $k^{(\text{TIE})}_{\text{opt}}-$based phase retrieval strategy, as a function of the resolution parameter $\mathcal{D}$. The results, shown in Fig. 7a, are reported for different defocus distances $dz$. Each data point represents the experimental resolution extracted from phase images averaged over $10^{3}$ frames to eliminate residual shot-noise. The shaded area indicates the theoretical resolution limit achievable with the current optical system. Fig. 7b displays representative phase reconstructions, also averaged over $10^{3}$ frames, for two defocus distances, $dz=0.025\,\text{mm}$ and $dz=0.1\,\text{mm}$, at three different values of $\mathcal{D}$. Rectangular regions within each phase image are magnified to illustrate resolution variations more clearly.
[…]
To evaluate the spatial resolution of our imaging system, we have analyzed the phase profiles extracted along a line perpendicular to the edge of the phase samples. These intensity profiles are fitted using an error function (Edge Spread Function, ESF) centered at position $x_{0}$:
@@
To quantify the uncertainty, we first determine the 95% confidence interval for the fitted parameter $w$, indicated as $w_{\text{sub}},w_{\text{sup}}]$. The standard error for $w$ is then approximated as $se_{w}=(w_{\text{sup}}-w_{\text{sub}})/3.92$ according to the $z$-test. Consequently, the standard error of the resolution $R$ is calculated as:
@@ V Author contributions
AP, GO and IRB devised the scheme of high resolution NIQPI. AP, GO and IRB developed the software for simulations and data analysis. The experiment and the data analysis have been carried out by AP with the help of IRB. IRB and AP wrote the paper with the contribution of all the Authors. The non-biological sample has been designed, fabricated and characterized by SS. IRB and MG are equally last co-Authors. IRB supervised the project together with MG who is the director of the Research Sector in which the work has been developed.
VI Conflict of interest
The authors declare no conflicts of interest.